Find the general values of trigonometric equation :-CosA +SinA =Cos2A + Sin2A . Ans:- 2npi or (2npi)/2+pi/6.

Grade:12th pass

1 Answers

Anomy
6 years ago
According to me......
 
 
CosA +SinA = Cos2A +Sin2A
CosA - Cos2A= Sin2A - Sin A
2Sin3A/2SinA/2 = 2cos3A/2SinA/2
Sin3A/2 = Cos3A/2
Tan3A/2=1
Tan3A/2 = Tanπ/4
3A/2 = nπ+π/4
A=2nπ/3+π/6  ...where n belongs to real no.